A centrifuge with radius 0.5 m is rotating to θ = 0.40t2 . At t = 2.0 s, what are the magnitudes of the (a) angular velocity, (b) linear velocity, and (c) radial/centripetal acceleration?
Answer #1

a)

angular velocity is given by

at t=2 s

W=0.8*2=1.6 rad/s

b)

Linear velocity is given by

V=RW =0.5*1.6 =0.8 m/s

c)

radial acceleration

aC=W2R=1.62*0.5 =1.28 m/s2